Could you please help me with this problem
Write the equation of a hyperbola given foci (-8,0)and (8,0) with vertices
( -4 square root 3,0) and ( 4 square root 3, 0)

Answer by mukhopadhyay(490)

Center is at (0,0); transverse axis is on the x-axis
Standard equation is:
x^2/a^2 - y^2/b^2 = 1
c=8 (distance between center and focus)
a^2=(4 sqrt3)^2=48 (distance between center and vertex)
c^2=a^2+b^2
=>b^2=64-48=16
The required equation of the hyperbola is:
x^2/48-y^2/16=1